The fifteen-year bond yields 5.4% and has a coupon of 7.4%. If this yield to maturity remains unchanged, what will be its price one year hence? Assume annual coupon payments and a face value of $100. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Price            $

b. What is the total return to an investor who held the bond over this year? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Total return            
Answer #1

Price of the bond=Coupon rate*Face Value/Yield to Maturity*(1-1/(1+Yield to Maturity)^t)+Face Value/(1+Yield to Maturity)^t

1.
=7.4%*100/5.4%*(1-1/(1+5.4%)^14)+100/(1+5.4%)^14
=119.300546825

2.
Total return=yield to maturity=5.40%
Proof:
Price of the bond now=7.4%*100/5.4%*(1-1/(1+5.4%)^15)+100/(1+5.4%)^15=120.209247462

Total return=(Coupon rate*Face Value+Price of the bond 1 year from now)/Price of the bond now-1=(100*7.4%+119.300546825)/120.209247462-1=5.40%